# Enrich the content of the image Using Context-Aware Copy Paste

### v1

@@ Abstract
Data augmentation remains a widely utilized technique in deep learning, particularly in tasks such as image classification, semantic segmentation, and object detection. Among them, Copy-Paste is a simple yet effective method and gain great attention recently. However, existing Copy-Paste often overlook contextual relevance between source and target images, resulting in inconsistencies in generated outputs. To address this challenge, we propose a context-aware approach that integrates Bidirectional Latent Information Propagation (BLIP) for content extraction from source images. By matching extracted content information with category information, our method ensures cohesive integration of target objects using Segment Anything Model (SAM) and You Only Look Once (YOLO). This approach eliminates the need for manual annotation, offering an automated and user-friendly solution. Experimental evaluations across diverse datasets demonstrate the effectiveness of our method in enhancing data diversity and generating high-quality pseudo-images across various computer vision tasks.
Introduction
@@
Deep Learning-based approach has become the major paradigm in many computer vision tasks, ranging from classification to segmentation. These approaches outperform traditional ones in terms of accuracy and generalization. However, the bottleneck of deep learning is the quality and quantity of the training dataset. To obtain a dataset, a large volume of images needs to be annotated, which is labour-intensive and time-consuming. In segmentation task, annotating a single image takes 1.5 hours(?). How to generate high-quality, highly realistic datasets has become a research hotspot in recent years.
@@
Figure 1: Comparison between the Copy-Paste method (first row) and CACP (second row). The former overlooks the contextual relevance between the base and target images, leading to disharmony.
@@
Previous data augmentation methods increase the diversity of images by applying operations such as flipping, rotating, and adding blur and noise. However, these techniques fail to enhance the content of images at the object level. To address this issue, the Copy-Paste method (?) was proposed. The idea is straightforward and intuitive: objects from target images are pasted onto source images at random positions, resulting in images with enriched content.
@@
However, the Copy-Paste method has several drawbacks:
@@
  • Context Neglect: The method overlooks the contextual relationship between the copied objects and the target images. For instance, a penguin is unlikely to appear in a desert, and a giraffe is unlikely to be found on a soccer field. These contextually inappropriate images contribute little to the practical significance of the augmented dataset.
@@
  • Dependency on Masks: The original Copy-Paste approach relies on publicly available image-mask pairs to generate new images, requiring extra effort to extend its application boundaries. This process is not adaptable to scenarios where masks are unavailable.
@@
To address the aforementioned issues, we propose a novel approach named Context-Aware Copy-Paste (CACP), leveraging large language models (LLMs). Our method integrates several NLP-based models to ensure contextual relevance between the source and target images. The main procedure is as follows: The BLIP model (?) is utilized to generate captions for the source images(the images to be augmented); Object365 (?), a dataset containing 365 distinct classes, serves as the target image set. For a given source image, we calculate a similarity score between its caption and the category names in Object365 using a BERT-based semantic similarity model. The category with the highest similarity score is selected, and an image from this category is randomly chosen as the target image; YOLO-obj365 is employed to detect objects in the target image. The bounding box of the detected object is then processed by SAM (?) to obtain a pixel-level mask;Finally, The object, guided by the mask, is pasted onto the source image.
@@
Our approach can be applied to several computer vision tasks, including classification, object detection, and segmentation. Without requiring extra manual annotation, the target gallery can be easily extended to adapt to specific tasks. Moreover, the proposed module can be integrated with other techniques to enhance model performance. Our contributions can be summarized as follows:
@@
  • We have demonstrated that robust segmentation results can be achieved by combining Grad-CAM and YOLO as prompt generators.
  • We evaluated our approach by conducting experiments across multiple tasks. The results indicate that our method outperforms previous data augmentation techniques.
Figure 2: Comparison between SAM segmentation results using different prompts. Single Bounding Box prompt(upper row)tend to get incomplete masks. Bounding Box with GradCAM points can generate more accurate and robust masks.
Related work
Copy-Paste
Crop-Paste has been widely used in semi-supervised scenarios for its efficiency.(??). Dovornik(?) first propose copy-paste approach for object detection task based on visual context, which boost the performance in VOC07(?) dataset.However, they only use voc2012 as their target gallery, which is hard to apply the proposed method to other specific scenarios.Besides, they use a CNN classifier to obtain context infomation, which is insufficient compared to our BLIP-based approach. Golnaz(?) first proposed Copy-Paste data augmentation method in Instance Segmentation. As described in their paper, They claim that simply pasting objects randomly is able to provide solid gain on top of baselines. Although their approach is easy to implement, the random-related approach generate images lose the concordance compared to real images. The distribution of co-occurrence of objects are ignored. Zhao(?) propose X-Paste, which leverage zero-shot recognition models CLIP to make the approach scalable. X-Paste provides impressive improvements over the CenterNet2. Viktor Olsson(?) propose ClassMix, which generate augmentations by mixing unlabeled samples by leveraging on the network’s predictions for respecting objects boundaries. We are inspired by their work to combine the vision language models with copy-paste to generate augmented images in a sufficient way.
vision language model and Zero-shot Segmentation
Vision-Language Pre-training (VLP) has recently made very encouraging breakthrough. The ability of zero-shot and image-text alignment makes it possible to assist the copy-paste pipeline.
CLIP (Contrastive Language-Image Pre-Training)(?) is a neural network trained on a variety of (image, text) pairs. It can be instructed in natural language to predict the most relevant text snippet, given an image, without directly optimizing for the task.
BLIP(?) is a new VLP framework which transfers flexibly to both vision-language understanding and generation tasks. It utilizes the noisy web data achieving SOTA results on several benchmarks, which also perform very well on zero-shot tasks.
SAM is a promptable segmentation model which is trained on the largest segmentation dataset so far(?). It can transfer zero-shot to new unseen distributions and tasks. The author claims that its zero-shot performance is competitive or even superior to prior fully supervised results. The only bottleneck of SAM is the prompt. The segmentation results may unstable or unintended given insufficient prompts. SAM is widely used to guide data augmentation. (?) introduce SAMAug, g, a novel visual point augmentation method for the SAM that enhances interactive image segmentation performance. However, the point they proposed lack semantic insight which is not efficient and meaningful.
Motivation
Although copy-paste data augmentation methods have significantly improved computer vision tasks, traditional crop-paste pipelines have two notable limitations. First, the crop-paste method is challenging to scale effectively. Second, the semantic gap between the source image and the target image remains unaddressed.
Scalability and Extendability
Previous copy-paste methods heavily rely on image-mask pairs to perform operations. However, preparing masks for images is costly and time-consuming, making it challenging to apply these methods in a generic manner. To address this issue, previous copy-paste approaches have resorted to using public datasets with pixel-level annotations such as VOC2017 and CamVid.
However, these datasets are limited in the number of categories they cover, thereby restricting the diversity of content in generated images. The table 1 provides descriptions of several public segmentation datasets (ADE20K (?), COCO (?), VOC2007 (?)). These datasets often cannot meet the specific requirements of scenarios. For instance, in a foreign object detection task, foreign objects may span hundreds of categories, making it impractical to rely on public datasets or manual annotation to prepare the dataset. An entirely automated crop-paste pipeline is needed to generate large quantities of high-quality images.
Dataset num Category resolution
VOC2007 9963 20 -
CamVid 701 32 480*360
CityScape 20000 30 2024*1048
coco 330k 80 640*480
ADE20K 25574 150 1650*2220
Table 1: Descriptions of public Segmentation datasets
Content dissonance
Previous work has often overlooked the issue of content relevance in the data augmentation process. Irrelevant objects are frequently pasted onto source images, providing minimal training benefit. As depicted in the figure below, images generated by traditional crop-paste methods contribute less to person-related computer vision tasks, as evidenced by corresponding Grad-CAM results.
We propose that incorporating highly relevant objects into source images at appropriate positions can enrich image content and expedite the training process. Experimental results validate our hypothesis: traditionally pasted parts fail to activate appropriately, while our method successfully triggers activations, enhancing the content relevant to desired classes.
Figure 3: GradCam comparison between the Copy-Paste(top row) and our context-aware copy paste(bottom row).
Method
In general, our CACP can be summarized in following parts: Gallery Preparation, Context-Aware, Crop Paste. In Gallery preparation part, target images are selected to provide object-level content enhancement; In context aware stage, source image and target image are bridged using BLIP and Bert-based similarity measurement tool. Once the preferred target image category is determined, object of interests in target image will be cropped and pasted onto the source image considering the size and position. The yolo and SAM will be leveraged to make the crop-paste process fully automatically. The details of each procedure will be described in follow paragraphs.
BLIP YOLO SAM total
Model Size(MB) 1824 14.2 2384 -
CPU(ms) 3687 97 3787 7571
GPU(ms) 1137 12 498 1647
Table 2: Model sizes and inference times in CPU and GPU
@@
Reformulation
@@
Given a source image sets $S_{S}$ and a target image $I_{T}$, our task is to find the most relevant image $I_{S}\in S_{S}$ and most relevant object $obj\in I_{S}$. Specifically, $I_{S}$ can be obtained as follow:
@@
where $I_{syn}$ is the generated image.
Data preparation
To enhance the diversity of our dataset, a substantial collection of images is essential for our gallery. In this study, we utilize Object365 (?) as our image gallery. Object365 encompasses 365 categories, featuring over 2 million images and 30 million bounding boxes. These images are characterized by high resolution and quality annotations. In contrast, COCO offers only 80 classes. Object365 significantly expands the range of target objects available for augmentation.
@@
Context-Aware
Image Caption
To establish semantic coherence between the source and target images, it is crucial to recognize the contents of both images beforehand. Rather than solely detecting objects within the images, we employ a state-of-the-art Visual-Language pre-training (VLM) model as the content extractor. In this role, we utilize BLIP (Bootstrapped Language-Image Pre-training). Compared to object-detection methods, BLIP generates smooth and natural descriptions of input images, rather than isolated words. Furthermore, while object-level approaches may struggle to provide meaningful information when encountering unseen objects, BLIP consistently offers general information applicable to common scenarios.
@@
In last step we obtain the caption of source image,namely $C(I_{src})$. Due to the large amount of the target image gallery, as a trade-off, we take the category name as the caption of the target image, annotated as $C(I_{T_{i}}(i=1,2,...,n))$,where n is the total number of categories. To determine the correlation between the $C(I_{src})$ and $C(I_{T_{i}})$, Bert-embedding is leveraged as our measurement tool. Below are samples using Bert-embedding to calculate the similarity in our work compared to traditional approaches.
@@
From the table we can find that Bert-based distance metric is preferred.
Crop-Paste
Once the category with the highest similarity score is determined, we randomly select an image from this category as the target image. The Segment Anything Model (SAM) is then employed as the pixel-level mask extractor. SAM is a single-shot segmentation model capable of segmenting any object based on prompts, such as bounding boxes or multiple points indicating the intended objects.
Prompts generation
To obtain these prompts, we utilize YOLO-365, a model trained on the Object365 dataset, to detect objects within the target images. For instance, if the YOLO-365 model detects a ’dog’ within a target image categorized under dogs, the bounding box of the dog is then forwarded to the SAM model along with the target image. SAM subsequently generates the corresponding segmentation mask for the detected object. Finally, guided by this generated mask, the pixels representing the dog are pasted onto the source images.
In our experimental trials, we observed that feeding pure bounding boxes into the Segment Anything Model (SAM) often results in unintended or incomplete masks for the corresponding target objects. To mitigate this issue, we propose an approach based on Grad-CAM (?) to achieve more accurate segmentation masks.
@@
Scale and Position
To enhance the realism and harmony of the generated image, rescaling and rendering techniques are implemented. The cropped objects are rescaled according to a ratio interval based on our statistical analysis of the Object365 dataset. We traversal the images in Objects365, and record ratios of image pair, namely ¡obj1, obj2, ratio¿. Before pasting the object onto the source images, we extract object pair names and obtain the $ratio_{max}^{obj1,obj2}$ and $ratio_{min}^{obj1,obj2}$ from record.
Experiments
@@
Car Pedestrian Building Road Sky Tree
U-Net(?) 0.776 0.447 0.764 0.872 0.872 0.834
U-Net+CACP 0.789 0.481 0.783 0.893 0.875 0.841
FPN(?) 0.792 0.432 0.783 0.897 0.884 0.867
FPN+CACP 0.813 0.479 0.797 0.903 0.885 0.881
PSPNet(?) 0.788 0.445 0.792 0.886 0.875 0.843
PSPNet+CACP 0.803 0.487 0.797 0.901 0.873 0.862
DeepLabV3(?) 0.792 0.461 0.803 0.908 0.891 0.875
DeepLabV3+CACP 0.811 0.493 0.812 0.922 0.887 0.893
DeepLabv3plus(?) 0.803 0.471 0.817 0.927 0.907 0.871
DeepLabv3plus+CACP 0.817 0.497 0.826 0.933 0.912 0.883
PAN(?) 0.794 0.501 0.806 0.931 0.883 0.868
PAN+CACP 0.812 0.513 0.821 0.937 0.891 0.891
Table 4: CACP works well across different segmentation architectures
Configuration
The experiment is conducted in pytorch platform. GPU is RTX 3090ti with 24GB memory. For classification task, Batch size is set to 16, loss function is cross entropy. Adam(?) is leveraged as optimizer, and learning rate is 0.001, epoch is set to 50; As for segmentation task, batch size is set to 8,loss function is dice loss,epoch is set to 20. As for detection task, we set batch size as 8, and epoch is set as 50. The models for segmentation task and object detection task we use the segmentation pytorch model and yolov5-s, respectively. Our segmentation models are implemented by segmentation models pytorch(?).
Metric
Precision is selected as the metric for classification task, which is calculated as below:
$Accuracy=\frac{Number\_correct\_predictions}{Totl\_number\_of\_Predictions}$ (4)
As for the segmentation task, mIoU is leveraged to evaluate the model’s performance.

### v2

# CACP: Context-Aware Copy-Paste to Enrich Image Content for Data Augmentation

@@ Abstract
Data augmentation is a widely used technique in deep learning, encompassing both pixel-level and object-level manipulations of images. Among these techniques, Copy-Paste stands out as a simple yet effective method. However, current Copy-Paste approaches either overlook the contextual relevance between source and target images, leading to inconsistencies in the generated outputs, or heavily depend on manual annotations, which limits their scalability for large-scale automated image generation. To address these limitations, we propose a context-aware approach that integrates Bidirectional Latent Information Propagation (BLIP) for extracting content from source images. By aligning the extracted content with category information, our method ensures coherent integration of target objects through the use of the Segment Anything Model (SAM) and YOLO. This approach eliminates the need for manual annotation, offering an automated and user-friendly solution. Experimental evaluations across various datasets and tasks demonstrate the effectiveness of our method in enhancing data diversity and generating high-quality pseudo-images for a wide range of computer vision applications.
1 Introduction
@@
Deep Learning-based approaches have become the major paradigm in many computer vision tasks, ranging from classification to segmentation. These approaches outperform traditional ones in terms of accuracy and generalization. However, the bottleneck of supervised deep learning is the quality and quantity of the training dataset. To obtain a dataset, a large volume of images needs to be annotated, which is labour-intensive and time-consuming. For segmentation task, annotating a single image is estimated to take up to 1.5 hours[7]. How to generate high-quality, highly realistic datasets has become an important research question in recent years.
@@
Figure 1: Comparison between the Copy-Paste method (first row) and CACP (second row). The former overlooks the contextual relevance between the base and target images, leading to disharmony. Our approach leverage the semantic information using CAM(Context-Aware-Module) to alleviate this issue.
@@
Previous data augmentation methods increase the diversity of images by applying operations such as flipping, rotating, and adding blur and noise. However, these techniques fail to enhance the content of images at the object level. To address this issue, the Copy-Paste method [14] was proposed. The idea is straightforward and intuitive: objects from target images are pasted onto source images at random positions, resulting in images with enriched content.
@@
However, existing Copy-Paste methods have several drawbacks:
@@
  • Context Neglect: Methods often neglect the contextual relationship between the copied objects and the target images. For example, a penguin is unlikely to appear in a desert, and a giraffe is improbable on a soccer field. Such contextually incompatible images reduce the practical significance of the augmented dataset.
@@
  • Dependency on Masks: The original Copy-Paste approach depends on publicly available image-mask pairs to generate new images, which limits its applicability and requires additional effort to extend its use. This process is not adaptable to scenarios where masks are unavailable.
@@
To address the aforementioned issues, we propose a novel approach named Context-Aware Copy-Paste (CACP), leveraging large language models (LLMs) and vision foundation modles. Our method integrates several NLP-based models to ensure contextual relevance between the source and target images. The main procedure is as follows: A vision-language model is used to generate captions for the source images (the images to be augmented). Object365, a dataset containing 365 distinct classes, serves as the target image set. For a given source image, a similarity score is computed between its caption and the category names in Object365 using a semantic similarity model based on a transformer architecture. The category with the highest similarity score is selected, and an image from this category is randomly chosen as the target image. An object detection model is then employed to identify objects in the target image. The bounding box of the detected object is processed by a segmentation model to obtain a pixel-level mask. Finally, the object, guided by the mask, is pasted onto the source image.
@@
Our approach can be applied to several computer vision tasks, including classification, object detection, and segmentation. Without requiring extra manual annotation, the target gallery can be easily extended to adapt to specific tasks. Our contributions can be summarized as follows:
@@
  • We demonstrate that robust segmentation results can be achieved by combining object detector and class activation mapping as prompt generators.
  • The experiments results demonstrate that our method outperforms the original Copy-Paste technique and enhances model performance.
@@
2 Related Work
Copy-Paste for Data Augmentation
Copy-Paste has been widely used in semi-supervised scenarios due to its efficiency [16, 15]. Dvornik et al.[12] first proposed the copy-paste approach for object detection tasks based on visual context, which significantly boosted performance on the VOC07 [13] dataset. However, they only used VOC2012[13] as their target gallery, making it challenging to apply the method to other specific scenarios. Additionally, they used a CNN classifier to obtain context information(describing an image only by a word), which is less effective compared to our BLIP-based approach(describing an image by a sentence). Golnaz et al.[14] was the first to propose the Copy-Paste data augmentation method for instance segmentation. They claim that simply pasting objects randomly provides substantial gains over baselines. Although their approach is easy to implement, the random pasting generates images that lack the grounding of real images, as the distribution of object co-occurrences is ignored. Zhao et al. [33] proposed X-Paste, which leverages zero-shot recognition models like CLIP to make the approach scalable. X-Paste demonstrated impressive improvements over CenterNet2. Viktor Olsson el al.[24] introduced ClassMix, which generates augmentations by mixing unlabeled samples based on the network’s predictions to respect object boundaries. Inspired by their work, we combined vision-language models with copy-paste to generate augmented images efficiently.
Vision Language Models and Zero-shot Segmentation
Vision-Language Pre-training (VLP) has recently made significant breakthroughs. The zero-shot capability and image-text alignment make it an ideal support for the copy-paste pipeline.
CLIP [26] is a neural network trained on a variety of (image-text) pairs. It can be instructed in natural language to predict the most relevant text snippet for a given image, without direct task-specific optimization. MaskCLIP[11]incorporates a newly proposed masked self-distillation into contrastive language-image pretraining, it distills representation from a full image to the representation predicted from a masked image. BLIP [22] is a new VLP framework that flexibly transfers to both vision-language understanding and generation tasks. Utilizing noisy web data, BLIP achieves state-of-the-art results on several benchmarks and performs exceptionally well on zero-shot tasks.
SAM is a promptable instance segmentation model trained on the largest segmentation dataset to date [20]. It can generalize to new, unseen distributions and tasks, with competitive or even superior performance compared to prior fully supervised results. However, SAM’s performance depends heavily on the quality of prompts; insufficient prompts can lead to unstable or unintended segmentation results. SAM is widely used to guide data augmentation. For example, [8] introduced SAMAug, a novel visual point augmentation method for SAM that enhances interactive image segmentation performance. The above approach could be improved by introducing semantic tool to bridge the source and target objects.
3 Motivation
Although copy-paste data augmentation methods have significantly improved computer vision tasks[14], traditional crop-paste pipelines have two notable limitations. First, the crop-paste method is challenging to scale effectively. Second, the semantic gap between the source image and the target image remains unaddressed.
3.1 Scalability and Extendability
Previous copy-paste methods heavily relied on image-mask pairs to perform operations. However, preparing masks for images is costly and time-consuming, making it challenging to apply these methods in a generic manner. To address this issue, previous copy-paste approaches have resorted to using public datasets with pixel-level annotations such as VOC2007 and CamVid[2, 13].
However, these datasets are limited in the number of categories they cover, thereby restricting the diversity of content in generated images. Tab. 1 provides a listing of properties of several public segmentation datasets (ADE20K [34], COCO [23], VOC2007 [13]). These datasets often cannot meet the specific requirements of scenarios. For instance, in a foreign object detection task[29], foreign objects may span hundreds of categories, making it impractical to rely on public datasets or manual annotation to prepare the dataset. An entirely automated copy-paste pipeline is needed to generate large quantities of high-quality images.
dataset images classes resolution
VOC2007[13] 9963 20 -
CamVid[1] 701 32 480*360
CityScape[6] 20000 30 2024*1048
coco[23] 330k 80 640*480
ADE20K[34] 25574 150 1650*2220
Table 1: Summary of properties of common public segmentation datasets
3.2 Content Discrepancy
Previous work has often overlooked the issue of content relevance in the data augmentation process. Irrelevant objects are frequently pasted onto source images, providing minimal training benefit. As depicted in Fig. 4 , images generated by traditional copy-paste methods contribute less to person-related computer vision tasks, as evidenced by corresponding Grad-CAM results. In addition, models trained on such images failed to learn scene context, i.e. which classes are likely to co-exist in an image.
We propose that incorporating highly relevant objects into source images at appropriate positions and scales can enrich the image content and expedite the training process. Experimental results validate our hypothesis: traditionally pasted elements often fail to activate appropriately. In contrast, our method successfully triggers activations, as demonstrated by GradCam visualization techniques Fig. 3, enhancing the content relevant to the desired classes.
Figure 3: GradCam comparison between the Copy-Paste(top row) and our context-aware copy paste(bottom row). CACP contributes more in person related vision tasks compared to copy-paste.
4 Method
Our CACP approach can be split into the following parts: Gallery Preparation, Context-Awareness Module and Copy-Paste. In the Gallery preparation part, target images are selected to provide object-level content enhancement; In the context aware stage, source image and target image are bridged using a BLIP and BERT-based similarity measurement tool. Once the preferred target image category is determined, object of interests in target image will be cropped and pasted onto the source image considering the size and position. An object detector model and SAM will be leveraged to make the crop-paste process fully automatic. The details of each procedure will be described in follow paragraphs.
Figure 4: Comparison of SAM segmentation results using different prompts: Single bounding box prompts (upper row) tend to produce incomplete masks, while combining bounding boxes with Grad-CAM points generates more accurate and robust masks.
4.1 Problem Formulation
@@
Given a source image set $\mathcal{D}_{S}$ and a target image $I_{T}$, our task is to find the most relevant image $I_{S}\in\mathcal{D}_{S}$ and most relevant object $obj\in\mathcal{O}$, where $\mathcal{O}$ is the collection of objects in $I_{S}$. Specifically, $I_{S}$ can be obtained as follow:
@@
where $I_{syn}$ is the generated image, $\otimes$ is pixel-wise multiplication.
4.2 Data preparation
To enhance the diversity of our dataset, a substantial collection of images is essential for our gallery. In this study, we utilize Object365 [29] as our image gallery. Object365 encompasses 365 classes, featuring over 2 million images and 30 million bounding boxes. These images are characterized by high resolution and quality annotations. In contrast, COCO offers only 80 classes. Object365 significantly expands the range of target objects available for augmentation.
@@ 4.2 Data preparation
4.3 Context Awareness Module
Image Captioning
To establish semantic coherence between the source and target images, it is crucial to recognize the contents of both images beforehand. Rather than solely detecting objects within the images, we employ a state-of-the-art Visual-Language pre-training (VLM) model as the content extractor. In this role, we utilize BLIP[22]. Compared to object-detection methods, BLIP generates smooth and natural descriptions of input images, rather than isolated words. Furthermore, while object-level approaches may struggle to provide meaningful information when encountering unseen objects, BLIP consistently offers general information applicable to common scenarios.
@@ 4.3 Context Awareness Module
In the last step we obtain the caption of the source image,namely $C(I_{S})$. Due to the large amount of the target image gallery, as a trade-off, we take the class name as the caption of the target image, annotated as $C(I_{T_{i}}(i=1,2,...,n))$,where n is the total number of categories. To determine the correlation between the $C(I_{S})$ and $C(I_{T_{i}})$, Bert-embedding[10] is leveraged as our measurement tool. In Table 2 we present examples of samples using Bert-embedding to calculate the similarity in our work compared to traditional approaches.
@@
From the Tab. 2 we can find that Bert-based distance metric is preferred.
4.4 Copy-Paste
Once the category with the highest similarity score is determined, we randomly select an image from this class as the target image. The SAM is then employed as the pixel-level mask extractor. SAM is a single-shot segmentation model capable of segmenting any object based on prompts, such as bounding boxes or multiple points indicating the intended objects.
4.4.1 Prompt generation
To obtain the prompts for segmentation, we utilize YOLO-365, a model trained on the Object365 dataset, to detect objects within the target images. For instance, if the YOLO-365 model detects a ’dog’ within a target image categorized under dogs, the bounding box of the dog is then forwarded to the SAM model along with the target image. SAM subsequently generates the corresponding segmentation mask for the detected object. Finally, guided by this generated mask, the pixels representing the dog are pasted onto the source images.
In our experimental trials, we observed that feeding pure bounding boxes into the Segment Anything Model (SAM) often results in unintended or incomplete masks for the corresponding target objects. To mitigate this issue, we propose an approach based on Grad-CAM [28] to achieve more accurate segmentation masks.
@@ 4.4.1 Prompt generation
4.4.2 Scale and Position
To enhance the realism and harmony of the generated image, rescaling and rendering techniques are implemented. The cropped objects are rescaled according to a ratio interval based on our statistical analysis of the Object365 dataset. We traverse the images in Objects365, and record ratios of image pair, namely $obj1-obj2-ratio$. Before pasting the object onto the source images, we extract object pair names and obtain the $ratio_{max}^{obj1,obj2}$ and $ratio_{min}^{obj1,obj2}$ from record.
5 Experiments
@@
Car Pedestrian Building Road Sky Tree $\uparrow$
U-Net[27] 0.776 0.447 0.764 0.872 0.872 0.834
U-Net+CACP 0.789(+0.013) 0.481(+0.034) 0.783+0.019 0.893(+0.021) 0.875(+0.003) 0.841(+0.007) +0.016
FPN[19] 0.792 0.432 0.783 0.897 0.884 0.867
FPN+CACP 0.813(+0.011) 0.479(+0.047) 0.797(+0.014) 0.903(+0.006) 0.885(+0.001) 0.881(+0.014) +0.015
PSPNet[32] 0.788 0.445 0.792 0.886 0.875 0.843
PSPNet+CACP 0.803(+0.015) 0.487(+0.013) 0.799(+0.007) 0.901(+0.015) 0.873(+0.002) 0.862(+0.019) +0.012
DeepLabV3[4] 0.792 0.461 0.803 0.908 0.891 0.875
DeepLabV3+CACP 0.811(+0.019) 0.493(+0.028) 0.812(+0.009) 0.922(+0.014) 0.887(-0.004) 0.89(+0.019) +0.011
DeepLabv3plus[5] 0.803 0.471 0.817 0.927 0.907 0.871
DeepLabv3plus+CACP 0.817(+0.014) 0.497(+0.026) 0.826(+0.009) 0.933(+0.006) 0.912(+0.005) 0.883(+0.012) +0.012
PAN[21] 0.794 0.501 0.806 0.931 0.883 0.868
PAN+CACP 0.812(+0.008) 0.513(+0.012) 0.821(+0.015) 0.937(+0.006) 0.891(+0.008) 0.891(+0.023) +0.012
Table 3: CACP provides robust gains across popular segmentation architectures in CamVid except Sky in DeeplabV3.
5.1 Configuration
The experiments are conducted in the pytorch platform. GPU is RTX 3090ti with 24GB memory. For the classification task, the batch size is set to 16 and the loss function is cross entropy. Adam[18] is used as the optimizer with learning rate of 0.001 over 50 epochs. For the segmentation task, batch size is set to 8, loss function is dice loss,epochs are set to 20. For the detection task, we set batch size as 8, and epochs are set to 50. For segmentation task and object detection task we use the segmentation pytorch model and yolov5-s[30], respectively. Our segmentation models are implemented by Segmentation-Models-Pytorch(SMP) library[17].
5.2 Metric
Precision is selected as the metric for classification task, which is calculated as given below:
$Accuracy=\frac{Number\_correct\_predictions}{Total\_number\_of\_Predictions}$ (4)
For the segmentation task, the mean Intersection over Union (mIoU) is used to evaluate the model’s performance, computed as follows: